What does this function return if it is called with n = 4?

int f(int n) {

if (n == 0)

return 1;

else

return f(n-1) + 1;

}

Can you show the steps as well please?
Answer #1

n = 4
to calculate f(4). it first calls f(3)
to calculate f(3). it first calls f(2)
to calculate f(2). it first calls f(1)
to calculate f(1). it first calls f(0)

so, let's go in bottom-up approach to solve this.
f(0) = 1    (base case)
f(1) = f(0) + 1 = 1 + 1 = 2
f(2) = f(1) + 1 = 2 + 1 = 3
f(3) = f(2) + 1 = 3 + 1 = 4
f(4) = f(3) + 1 = 4 + 1 = 5

so, f(4) gives a value of 5.

Answer: 5